OpenStudy (anonymous):

Of all the Harry Potter books purchased in a recent year, about 60% were purchased for readers 14 or older. If 12 Harry Potter fans who bought books that year are surveyed, find the following probabilities At least five of them are 14 or older. Exactly nine of them are 14 or older Less than three of them are 14 or older.

terenzreignz (terenzreignz):

Ideas? It amounts to saying that each unspecified reader has a 60% (or 3 out of 5) chance of being 14 or older, right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so do i take 60% of 12?

OpenStudy (anonymous):

how would i solve the other two problems

terenzreignz (terenzreignz):

Hold on there XD It's not that simple...

OpenStudy (anonymous):

could you explain then? Beacuse I really don't get this question

OpenStudy (anonymous):

because

terenzreignz (terenzreignz):

Independent events, heard of them?

OpenStudy (anonymous):

yeah

terenzreignz (terenzreignz):

what are they?

OpenStudy (anonymous):

the occurrence of one event doesn't effect the chance that the other will occur or not

terenzreignz (terenzreignz):

Good. So, the status of one of the twelve readers being a 14-and-above doesn't affect the others' right?

OpenStudy (anonymous):

yeah

terenzreignz (terenzreignz):

One thing about independent events, if you want all of them to happen at the same time, you just multiply probabilities :) Now, the probability for ONE of the readers to be a 14-and-older is 60% or 0.6, yes?

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

The probability for at least two of them to be 14-and-older, being independent, you simply multiply the two... 0.6*0.6 = 0.36 or 36% But you need at least FIVE. Work it out :)

OpenStudy (anonymous):

7.76%

OpenStudy (anonymous):

sorry 7.78%

terenzreignz (terenzreignz):

all right, hang on...

terenzreignz (terenzreignz):

Wait, I think I screwed up...

OpenStudy (anonymous):

oh ok

terenzreignz (terenzreignz):

Is this binomial distribution?

OpenStudy (anonymous):

i think

terenzreignz (terenzreignz):

Even I oversimplified it LOL Okay, clean slate

OpenStudy (anonymous):

ok

terenzreignz (terenzreignz):

Formula for binomial distribution, do you know it?

OpenStudy (anonymous):

yeah

terenzreignz (terenzreignz):

State it.

terenzreignz (terenzreignz):

Okay, let's let p be the probability of the reader being 14 or older. p = 0.6, right?

OpenStudy (anonymous):

yeah

terenzreignz (terenzreignz):

Now, the probability of there being EXACTLY r readers in the 12 being 14 or older is given by \[\Large \left(\begin{matrix}12\\r\end{matrix} \right)p^r(1-p)^{12-r}\]

terenzreignz (terenzreignz):

With this, you can answer the third question immediately. What's the answer to the third?

terenzreignz (terenzreignz):

Sorry, I meant the second one...

OpenStudy (anonymous):

0.142

terenzreignz (terenzreignz):

Good. So much for the second. For the first one, we'll have to add. So, when you want the probability for 'at least 5', you might have to add "exactly 5" "exactly 6" ... and so on, until "exactly 12" Or...

terenzreignz (terenzreignz):

There is a slightly faster way :)

terenzreignz (terenzreignz):

Complements. Do you know them?

OpenStudy (anonymous):

no. so for the first one i just have to add 0.6 five times?

terenzreignz (terenzreignz):

No. You do something similar to what you did in the second question (replace r with 9), but you'll do it like...eight times (you'll have to find the probabilities from 5 to 12) Unless... you use complements.

OpenStudy (anonymous):

so i have to do that same equation 8 times....from 9, 8, 7, 6, etc for the values of r and then add them?

terenzreignz (terenzreignz):

What you do is instead, count the instances of the event NOT happening. So, to negate "at least 5" we need "less than 5" So check the probability of... -no reader 14 and above (r = 0) -one reader (r = 1) -two readers (r = 2) -three readers (r= 3) -four readers (r = 4) And then add them up and tell me what you get.

OpenStudy (anonymous):

so for the third one do i just do the same thing as the first and then add them up?

terenzreignz (terenzreignz):

First one's not done yet. After you add up (this is the probability of the desired event NOT happening, remember?) Subtract it from 1, and that'll be your answer to the first. for the third, add up the probabilities involving things less than 3 (r = 0,1,2)

OpenStudy (anonymous):

and then i will be done?